A circular disc of mass 5kg and diameter 0.2m initially at rest attains a angular velocity of 600rpm in 10 sec. Calculate the torque acting on it and the rate at which work is being done at the end of 10 sec
1
Expert's answer
2013-01-30T07:44:06-0500
Moment of inertia I = 0.5*m*r^2
torque T = I*acc

angular acceleration acc = w / t

torque T = 0.5*m*r^2 * w / t = 0.5* 5 kg *(0.2 m)^2 * (600 1/min) / (10 sec) =
0.1 N*m

rate at which work is being done is power P = T * w = 0.1 N*m * (600 1/min) = 1
W

Answer: torque 0.1 N*m, power 1 W